A belt pulley <b><i>system</i></b> uses a flat belt of cross section area 800mm squared and density <b><i>1200 kg</i></b>/m cubed. Angle of lap is 160 degrees on the smaller pulley. Coefficient of friction is 0.3. Maximum stress allowed in the belt is 3N/mm squared. Calculate; (i) Maximum power transmitted by varying speed. (ii) Speed at which it occurs. (iii The initial tension of the belt


1
Expert's answer
2021-07-27T02:38:01-0400

Part i

"P_{max}=σ_b⋅A⋅k⋅v\\
P_{max}=3*800*0.3⋅0.2887*10^3*\frac{160}{360}\\
P_{max}=92384 W\\
P_{max}= 92.384 kW"


Part ii

"v= \sqrt{\frac{\sigma_p-\sigma_b}{3 \rho}}\\
v= \sqrt{\frac{(3-0)*10^3}{3 * 12000}}\\
v=0.2887 m/s"

"v= \frac{\pi d_1N}{60}\\
v= \frac{\pi \sqrt\frac{4A}{\pi}N}{60}\\
0.2887*10^3= \frac{\pi \sqrt\frac{4*800}{\pi}N}{60}\\
N = \frac{0.2887*10^3}{ \frac{\pi \sqrt\frac{4*800}{\pi}}{60}}\\
N= 172.76 rpm"


Part iii

"\frac{T_1}{T_2}=e^{\mu \theta}\\
\frac{T_1}{T_2}=e^{0.3* \frac{160}{360}\pi}\\
\frac{T_1}{T_2}=1.52\\
Also, \\
3000=\frac{T_1+T_2}{2}\\
\therefore T_1 = 1.52T_2= 1.52(6000-T_1)\\
T_1 = 9120-1.52T_1\\
T_1 = \frac{9120}{2.52}\\
T_1=3619.05 N"